fatimah is 5 years older than dollah.if the product of their ages is 234.how old is dollah?
Answer by Cromlix(4381)

Hi there,
Make Dollah's age = x
Fatimah's age = x + 5
The product of their ages = 234
x(x + 5) = 234
Multiply out
x^2 + 5x = 234
Sort into ax^2 + bx + c = 0 form
x^2 + 5x - 234 = 0
(x - 13)(x + 18) = 0
x + 18 = 0
x = -18 (Disregard as -ve)
x - 13 = 0
x = 13
Dollah is 13 years old
Fatimah is 18 years old.
Hope this helps :-)